# Exciton self-trapping in twisted hexagonal boron nitride homostructures

### v1

@@
Sébastien Roux Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France sroux@insa-toulouse.fr    Christophe Arnold Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France    Etienne Carré Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France    Alexandre Plaud Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France    Lei Ren Université de Toulouse, INSA-CNRS-UPS, LPCNO, 135 Av. Rangueil, 31077 Toulouse, France    Eli Janzen Tim Taylor Department of Chemical Engineering, Kansas State University Manhattan, KS 66506, USA    James H. Edgar Tim Taylor Department of Chemical Engineering, Kansas State University Manhattan, KS 66506, USA    Camille Maestre Laboratoire des Multimatériaux et Interfaces, UMR CNRS 5615, Univ Lyon
@@
   Vincent Garnier Laboratoire MATEIS, UMR CNRS 5510, Univ Lyon, INSA Lyon, F-69621 Villeurbanne, France    Philippe Steyer Laboratoire MATEIS, UMR CNRS 5510, Univ Lyon, INSA Lyon, F-69621 Villeurbanne, France    Takashi Taniguchi Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan    Kenji Watanabe Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan    Cédric Robert Université de Toulouse, INSA-CNRS-UPS, LPCNO, 135 Av. Rangueil, 31077 Toulouse, France    Xavier Marie Université de Toulouse, INSA-CNRS-UPS, LPCNO, 135 Av. Rangueil, 31077 Toulouse, France    Annick Loiseau Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France    Julien Barjon Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France
(15 mai 2024)
@@
One of the main interests of 2D materials is their ability to be assembled with many degrees of freedom for tuning and manipulating excitonic properties. There is a need to understand how the structure of the interfaces between atomic layers influences exciton properties. Here we use cathodoluminescence (CL) and time-resolved CL experiments to study how excitons interact with the interface between two twisted hexagonal boron nitride (hBN) crystals with various angles. An efficient capture of free excitons by the interface is demonstrated, which leads to a population of long-lived and interface-localized (2D) excitons. Temperature dependent experiments indicate that for high twist angles, these excitons localized at the interface further undergo a self-trapping. It consists in a distortion of the lattice around the exciton on which the exciton traps itself. Our results suggest that this exciton-interface interaction causes a broad optical emission of highly twisted hBN-hBN structures around 300 nm (4 eV). Exciton self-trapping is finally discussed as a common feature of sp2 hybridized boron nitride polytypes and nanostructures due to the ionic nature of the B-N bond and their compact excitons.
@@
The conception of 2D material heterostructures (h2D) benefits from a large number of degrees of freedom in the choice of atomic layers and the way they are stacked, creating structural singularities at the interfaces between the layers [1, 2, 3, 4]. Taking advantage of these capabilities allows to control and manipulate the properties of excitons: the efficiency of their radiative recombination [5, 6, 7, 8, 9, 10], the properties of their emission [11], their diffusion length and their valley and/or spin coherence [12, 13], the dielectric screening between the electron and the hole [14, 15, 16, 17]. Finally, excitons can interact with Moiré-superpotentials, which tune their properties and motion [18, 19, 20, 21, 22, 23, 24]. Thus, h2Ds offer an ideal platform for creating novel electronic devices using exciton fluxes (excitronics) [25], or using their spin/valley indices (valleytronics and spintronics) [26]. Hexagonal boron nitride (hBN) is present in most h2D since it is the best insulating material for use as a substrate or as a capping layer for other 2D materials such as transition metal dichalcogenides (TMDs), 2D magnets [27] and graphene [28, 29, 11, 30, 31, 32, 14]. Because hBN plays such a key role in h2D-based devices, it is crucial to understand the surface and interface effects associated with it.
@@
The simplest hBN-based h2D is the twisted hBN-hBN homostructure, which consists of two hBN crystals with different in-plane crystal orientations. These structures exhibit many properties that are not present in single hBN crystals. At small twist angles, triangular ferroelectric domains appear at the interface [33, 34], as well as drastic changes in the electronic properties, such as in some cases, a spatial separation of the electron and hole wave functions within the Moiré supercell [35, 36]. At high twist angles, hBN-hBN homostructure exhibits intense second harmonic generation, which could be modulated by the twist angle [37]. Finally, a new luminescence signal has recently been discovered in cathodoluminescence (CL) on twisted hBN-hBN structures. It is characterized by a large linewidth (2 eV) and a maximum intensity at 4 eV (300 nm wavelength), i.e. 2 eV below the hBN gap [38, 39]. The origin of this optical emission and of the 2 eV redshift is debated. First, a giant exciton trapping at the interface Moiré was proposed [39, 40]. However, this is not consistent with theoretical studies, which estimate the depth of the interface energy well to be only a few hundreds of meV for excitons [41, 42]. Therefore, another group has discarded its excitonic nature and rather attributed this emission to deep defects near the interface [41].
@@
An alternative explanation involved the occurrence of exciton self-trapping at the interface of the homostructure. The phenomenon of self-trapping of an electron, hole, or exciton was predicted theoretically in the 1930s [43, 44] and demonstrated experimentally in alkali halide crystals as early as the 1960s [45, 46, 47, 48, 49]. Self-trapping results from a local deformation of the crystal lattice around a charged particle (or dipole) where the particle (or dipole) is trapped. The particle (or dipole) induces the lattice deformation, hence the term ”self-trapped state”. This state is a kind of polaron : a charge associated with the deformation cloud it induces around itself. If the coupling with the deformation mode is strong enough, the self-trapped state is more energetically stable than the Bloch state of the free particle in the undeformed lattice [43]. hBN is a good candidate for exciton self-trapping since it is a highly ionic material with a very compact exciton [50], which makes excitons and atomic B-N bonds dipoles of similar size, which is expected to favor exciton-lattice interactions.
@@
Today, the puzzle for understanding the nature and the origin of this new optical emission is incomplete. In this work, we present a deterministic approach to elucidate the interplay between twist angles, defects and excitons using continuous and time-resolved cathodoluminescence. To this end, sixteen hBN-hBN structures with different twist angles and fabricated from different hBN crystal sources are studied in cathodoluminescence (CL), in time-resolved CL (TRCL), at room temperature and as a function of temperature. We show how the whole set of experimental data support an exciton self-trapping mechanism occurring at the interface of hBN-hBN homostructures.
@@
All cathodoluminescence (CL) experiments were performed using a JEOL7001F scanning electron microscope (SEM) equipped with a Horiba Jobin-Yvon CL system optimized for UV detection, as described in detail in reference [51]. Secondary electrons are detected by a Everhart-Thornley detector (SE detector). The SEM image corresponds to the intensity of the signal collected by the SE detector as a function of the position of the focused electron beam, with an ultimate spatial resolution of 1.2 nm. Cathodoluminescence images are measured by a photomultiplier (Hamamatsu HJY model - cooled to -30°C by Peltier effect) as the electron beam scans the selected area of the sample. The signal from the SE detector and the intensity collected by the photomultiplier are recorded simultaneously. In this way, SEM and CL images of the same area are generated, which allows to correlate the emission of a CL signal with the topography of the sample. The CL image can be panchromatic (total intensity, unfiltered) or monochromatic, filtered in wavelength by monochromator equiped with diffraction gratings. Cathodoluminescence spectra are recorded with a CCD camera.
@@
Thanks to a careful calibration in intensity [51], the internal quantum efficiency ($IQE$) of a luminescence signal can be evaluated from a CL spectrum. It corresponds to the ratio between the rate of photons emitted inside the sample and the rate of electron-hole pairs generated by the incident electron beam. Details of the $IQE$ measurement for the considered signal are given in the Appendix A. Finally, time-resolved CL (TRCL) experiments are performed using a custom-built beam-blanking device mounted on the SEM column with a time resolution of 100 ps (details of the setup are presented in ref. [50]).
@@
To study the emission of twisted hBN-hBN structures, we decided to vary a large number of parameters in the fabrication of the homostructures: the twist angle, the thickness of the crystal flakes in stack, but also the quality of the starting hBN crystals and the way the two crystal flakes are assembled. hBN crystals with natural boron isotope content were grown by three different processes: an Atmospheric Pressure High Temperature (APHT) process [54, Hoffman2014, 52, 53] from Ni/Cr solvent, a High-Pressure High Temperature (HPHT) route [55, 56], and a Polymer Derived Ceramic (PDC) method [57, Maestre2022]. hBN from the HPHT method is recognized in the 2D materials scientific community as the reference hBN crystals. In a previous study, a quantitative benchmarking of the respective quality, in terms of defect density, of these different hBN sources was performed based on the measured free exciton lifetimes [50]. The values of the lifetime and $IQE$ of the free exciton measured on bulk crystals from the different hBN sources are given in Tab. 1.
To fabricate twisted homostructures made of HPHT crystals, the bottom hBN layer is mechanically exfoliated using a PDMS viscoelastic stamp and then deposited onto a Si/SiO2 substrate [59]. The second hBN flake is exfoliated in the same manner, aligned under a microscope over the first crystal flake and then deposited on top of it. Finally, the samples are annealed at 150°C for 1h30 on a hot plate. Note that with this method, one atomic layer that composes the twisted interface was in contact with the polymer.
For twisted homostructures made of APHT and PDC crystals, the bottom layer of hBN is mechanically peeled off using Scotch tape on a Si/SiO2 substrate, according to the method originally developed for graphene [60]. Since the interaction between the tape and the hBN flake is stronger than the Van der Waals interactions between the hBN layers, only part of the crystal can be detached and transferred on the substrate. The lower hBN surface is thus freshly cleaved. The second flake of hBN is peeled off using Scotch tape on PDMS, aligned, and then deposited onto the bottom hBN flake [59]. This method ensures that the two atomic layers that make up the hBN-hBN interface have never been in contact with either the PDMS or the Scotch tape. The procedures were performed in a controlled atmosphere glove box.
The twist angle between the two hBN flakes is measured in diffraction by electron backscatter diffraction (EBSD) or by backscattered electron channeling pattern (ECP), as shown in Fig. 1 for 3 HPHT samples. Details of the twist angle measurement are presented in the Appendix B. The samples are named according to the crystal synthesis method from which they are issued (HPHT, APHT or PDC) and the twist angle. Tab. 1 lists the different samples and the respective thickness of the top and bottom parts measured by Atomic Force Microscopy (AFM).
Sample 215 nm $IQE$ $\tau$ (ns) Top hBN Bottom hBN
(%) [50] [50] (nm) (nm)
@@
.3 III - Evidence of the excitonic origin of the broad 300 nm optical emission
@@
All hBN-hBN samples were first analyzed using continuous excitation CL (cw-CL) at room temperature. The first observation is that the luminescence at 300 nm mainly appears in homostructures with high twist angles (Fig. 2). At low twist, such as the PDC-2°, APHT-3° and APHT-4° samples, no luminescent signal could be detected in the 250-400 nm range, confirming a recent study [39]. We conclude that luminescence at 300 nm becomes truly significant at angles greater than 5°. A spectrum measured on a low angle hBN-hBN structure is presented in the Appendix C. Spectra obtained on 3 highly twisted hBN-hBN homostructures (purple) and on a single hBN crystals as a reference (black) are shown in Fig. 2(b). The broad 300 nm emission is modulated by Fabry-Pérot interferences whose periodicity is consistent with the flakes thickness (see Tab. 1). The signals are comparable for all samples, whatever the hBN source and flakes thicknesses. It displays common characteristics to previous observations [39, 40, 41]: very broad luminescence ranging from 250 to 400 nm (2 eV width), intense and centered around 300 nm, redshifted by 2 eV compared to the bulk hBN free exciton luminescence at 215 nm.
CL images filtered at 300 nm are presented in Fig. 2(c) and reveal that this emission is only detectable when the two twisted hBN crystals are in contact. These images clearly show that the 300 nm cathodoluminescence is associated with the presence of an interface between the two twisted hBN crystals. In summary of these first observations, the detection of the broad 300 nm optical emission appears independent of the quality of the hBN crystal, which varies greatly depending on the hBN synthesis process [50]. The emission phenomenon is therefore very robust. It appears unconditionally as soon as the twist angle is large, as already reported in Ref. [39] for instance.. As it does not depend on the differences in nature and density of defects that present the three hBN sources used to fabricate the homostructures, the emission seems to be intrinsic to the presence of the interface itself rather than related to extrinsic defects. This raises the question of its origin and more specifically of the interplay between the exciton and the interface of the homostructure, that we will explore in the following using TRCL.
@@
Figure 2: (a) SEM images of the twisted homostructures, outlined with a dotted line; (b) CL spectra measured on the bottom crystal alone (in black) and on the homostructure (in purple) at 300 K; (c) CL images filtered at 300 $\pm$ 7.5 nm at 300 K, on HPHT-11° (5 kV, 27 pA), APHT-29° (3 kV, 280 pA), and PDC-18° (5 kV, 340 pA) samples. Scale bars are 10 µm.
@@
The goal of the TRCL study is to identify the role of bulk free excitons in the process leading to the luminescence at 300 nm. It has been performed on HPHT based homostructures at room temperature. Fig. 3(a) compares the decay of the free exciton luminescence at 215 nm in the HPHT-11° homostructure with the decay obtained on the bottom hBN crystal alone. A drastic change is observed. For the crystal alone, the decay is dominated by a short 3 ns component, typical of the TRCL decay of the free exciton in a single bulk HPHT-hBN crystal [50]. For the homostructure, the fast initial component is similar but a long component of 0.46 µs appears with a significant intensity (30% on the HPHT-11° sample). This long time is much larger than the radiative lifetime of the free exciton measured at 27 ns [50]. This long-lived free exciton dynamics appear only in the presence of a twisted hBN-hBN interface. This could therefore be attributed to a significant detrapping of excitons from the interface. It shows the existence of a reservoir of excitons trapped at the interface and implies their prior trapping. TRCL thus reveals a phenomenon of trapping-detrapping of free excitons at the interface. It results in the coupling of the population of free excitons in the volume of the hBN crystals with the population of long-lived excitons captured at the interface between the twisted crystals.
@@
Figure 3: (a) Decay of the free exciton luminescence, filtered at 215 $\pm$ 7.5 nm, on the bottom hBN alone (gray) and on the hBN-hBN twisted structure HPHT-11° (orange), after interrupting the TRCL excitation at initial time (5 kV, 27 pA, spot size of 10 µm). Bi-exponential decay functions are used for the fits. (b) Decay of the free exciton luminescence (black), and of the the 300 nm emission (purple), filtered at 300 $\pm$ 7.5 nm, after interruption of steady excitation at t=0, on HPHT-11° homostructure (5 kV, 27 pA, 10 µm excitation spot diameter). Intensities during steady state excitation (t $<$ 0) are normalized by the integrated intensity of the bands recorded on continuous CL spectra. Experiments done at 300 K.
As shown in the Appendix D, the temporal decay of the broad 300 nm luminescence is independent of wavelength from 250 to 400 nm. This suggests that this large signal does not originate from an ensemble of quantum emitters with different energies.
Fig. 3(b) compares the decay of the free excitons in the hBN bulk crystal at 215 nm and the decay of the 300 nm luminescence both recorded in TRCL on the HPHT-11° twisted homostructure. Without going into the quantitative analysis of the characteristic times of the decays (presented later), we already note that the characteristic decay time of the the 300 nm luminescence is equal to the characteristic time of the long component of the free exciton at 215 nm. This means that the 300 nm luminescence is fed by the population of excitons trapped at the interface. The TRCL experiments demonstrate that the 300 nm luminescence originates from the trapping of free excitons at the interface. Still, the radiative process has to be identified, which is the subject of the next parts.
.4 IV - Power dependence and internal quantum efficiency of the 300 nm emission.
@@
During CL experiments, the electron beam excitation is spread in depth, so that the hBN-hBN interface is indirectly excited by the diffusion of free excitons generated in the volume of the hBN crystals towards the interface between them [61]. The efficiency of the interface exciton luminescence at 300 nm is then equal to the efficiency of the transport of free excitons from the volume to the interface times the efficiency of the radiative recombination process at the interface.
@@
The luminescence from the twisted hBN-hBN homostructures is analyzed here as a function of the areal density of the excitation power (W.cm-2). Low power density data were recorded with an acceleration voltage of 5 kV with a constant current of 27 pA and spot diameters ranging from 1.6 to 26 µm. High power densities were obtained with a fixed spot diameter of 1.6 µm and excitation currents ranging from 27 to 5 600 pA. This dual approach made it possible to vary the excitation power density over 5 orders of magnitude.

### v2

@@
Sébastien Roux Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France sroux@insa-toulouse.fr    Christophe Arnold Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France    Etienne Carré Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France    Alexandre Plaud Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France    Lei Ren Université de Toulouse, INSA-CNRS-UPS, LPCNO, 135 Av. Rangueil, 31077 Toulouse, France    Frédéric Fossard Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France    Nicolas Horezan Université Paris-Saclay, ONERA, DMAS, 92320 Châtillon, France    Eli Janzen Tim Taylor Department of Chemical Engineering, Kansas State University Manhattan, KS 66506, USA    James H. Edgar Tim Taylor Department of Chemical Engineering, Kansas State University Manhattan, KS 66506, USA    Camille Maestre Laboratoire des Multimatériaux et Interfaces, UMR CNRS 5615, Univ Lyon
@@
   Vincent Garnier Laboratoire MATEIS, UMR CNRS 5510, Univ Lyon, INSA Lyon, F-69621 Villeurbanne, France    Philippe Steyer Laboratoire MATEIS, UMR CNRS 5510, Univ Lyon, INSA Lyon, F-69621 Villeurbanne, France    Takashi Taniguchi Research Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan    Kenji Watanabe Research Center for Electronic and Optical Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan    Cédric Robert Université de Toulouse, INSA-CNRS-UPS, LPCNO, 135 Av. Rangueil, 31077 Toulouse, France    Xavier Marie Université de Toulouse, INSA-CNRS-UPS, LPCNO, 135 Av. Rangueil, 31077 Toulouse, France    François Ducastelle Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France    Annick Loiseau Université Paris-Saclay, ONERA-CNRS, LEM, 92320 Châtillon, France    Julien Barjon Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France
(27 février 2025)
@@
One of the main interests of 2D materials is their ability to be assembled with many degrees of freedom for tuning and manipulating excitonic properties. There is a need to understand how the structure of the interfaces between atomic layers influences exciton properties. Here we use cathodoluminescence and time-resolved cathodoluminescence experiments to study how excitons interact with the interface between two twisted hexagonal boron nitride (hBN) crystals with various angles. An efficient capture of free excitons by the interface is demonstrated, which leads to a population of long-lived and interface-localized (2D) excitons. Temperature dependent experiments indicate that for high twist angles, these excitons localized at the interface further undergo a self-trapping. It consists in a distortion of the lattice around the exciton on which the exciton traps itself. Our results suggest that this exciton-interface interaction causes the broad 4-eV optical emission of highly twisted hBN-hBN structures. Exciton self-trapping is finally discussed as a common feature of sp2 hybridized boron nitride polytypes and nanostructures due to the ionic nature of the B-N bond and the small size of their excitons.
@@
The conception of 2D material heterostructures (h2D) benefits from a large number of degrees of freedom in the choice of atomic layers and the way they are stacked, creating structural singularities at the interfaces between the layers [1, 2, 3, 4]. Taking advantage of these capabilities allows to control and manipulate the properties of excitons: the efficiency of their radiative recombination [5, 6, 7, 8, 9, 10], the properties of their emission [11], their diffusion length, their valley and/or spin coherence [12, 13] and the dielectric screening between the electron and the hole [14, 15, 16, 17]. Finally, excitons can interact with Moiré-superpotentials, which tune their properties and motion [18, 19, 20, 21, 22, 23, 24]. Thus, h2Ds offer an ideal platform for creating novel electronic devices using exciton fluxes (excitronics) [25], or using their spin/valley indices (valleytronics and spintronics) [26]. Hexagonal boron nitride (hBN) is present in most h2D since it is the best insulating material for use as a substrate or as a capping layer for other 2D materials such as transition metal dichalcogenides (TMDs), 2D magnets [27] and graphene [28, 29, 11, 30, 31, 32, 14]. Because hBN plays such a key role in h2D-based devices, it is crucial to understand the surface and interface effects associated with it.
@@
The simplest hBN-based h2D is the twisted hBN-hBN homostructure, which consists of two hBN crystals with different in-plane crystal orientations. These structures exhibit many properties that are not present in single hBN crystals. At small twist angles, triangular ferroelectric domains appear at the interface [33, 34], as well as drastic changes in the electronic properties, such as in some cases, a spatial separation of the electron and hole wave functions within the Moiré supercell [35, 36]. At high twist angles, hBN-hBN homostructure exhibits intense second harmonic generation, which could be modulated by the twist angle [37]. Finally, a new 4-eV luminescence signal has recently been discovered in cathodoluminescence (CL) on twisted hBN-hBN structures.
Twisted hBN-hBN structures are composed of two stacked multilayer hBN flakes, as shown in Fig. 1 (a). On the one hand, the single flakes are arranged in the AA’ stacking configuration (Fig. 1 (b)), with high symmetry and exclusively out-of-plane hetero $\pi^{B-N}$ bonds. On the other hand, the twisted interface presents a variety of out-of-plane bonds, including homo $\pi^{B-B}$ and $\pi^{N-N}$ bonds (Fig. 1 (c-d)). For small twist angles, a macroscopic moiré superstructure appears, with $\pi^{B-B}$ and $\pi^{N-N}$ localized at specific positions of the interface. For large twist angles, close to the 30° twist quasi-crystal limit, atomic orbitals vary locally from site to site, and homo $\pi^{B-B}$ and $\pi^{N-N}$ bonds are densely distributed all over the interface. The luminescence of single flakes is dominated by the free exciton emission, while highly twisted hBN-hBN structures exhibit a new emission, shown in Fig. 1 (e), characterized by a large linewidth (2 eV) and a maximum intensity at 4 eV, i.e. 2 eV below the hBN gap [38, 39]. The origin of this optical emission, its 2-eV redshift, and its relation to the structure of the interface are debated in the literature. First, a giant exciton trapping at the interface moiré was proposed [39, 40]. However, this is not consistent with theoretical studies, which estimate the depth of the interface energy well to be only a few hundreds of meV for excitons [41, 42]. Therefore, another group has discarded its excitonic nature and rather attributed this emission to deep defects near the interface [41].
Figure 1: (a) Sketch of a hBN-hBN structure. Top views of (b) the AA’ stacking arrangement of hBN, (c) a 3°-twisted hBN-hBN interface and (d) a 23°-twisted hBN-hBN interface. (e) CL spectra measured on a single bulk hBN crystal (the bottom part alone), in black, and on the 23° hBN-hBN stack, in red. The luminescence is modulated by Fabry-Perot interferences related to the thinness of the crystals.
@@
In this paper we propose an alternative explanation with the occurrence of exciton self-trapping at the interface of the homostructure. The phenomenon of self-trapping of an electron, hole, or exciton was predicted theoretically in the 1930s [43, 44] and demonstrated experimentally in alkali halide crystals as early as the 1960s [45, 46, 47, 48, 49]. Self-trapping results from a local deformation of the crystal lattice around a charged particle (or dipole) where the particle (or dipole) is trapped. The particle (or dipole) induces the lattice deformation, hence the term ”self-trapped state”. This state is a kind of polaron: a charge associated with the deformation cloud it induces around itself. If the coupling with the deformation mode is strong enough, the self-trapped state is more energetically stable than the Bloch state of the free particle in the undeformed lattice [43]. hBN is a good candidate for exciton self-trapping since it is a highly ionic material with a very compact exciton [50]. It makes excitons and atomic B-N bonds dipoles of similar size, which is expected to favor exciton-lattice interactions.
@@
Today, the puzzle for understanding the nature and the origin of this new optical emission is incomplete. In this work, we present a deterministic approach to elucidate the interplay between twist angles, defects and excitons using CL and time-resolved CL (TRCL). To this end, sixteen hBN-hBN structures with different twist angles and fabricated from different hBN crystal sources are studied in CL and TRCL, at room temperature and as a function of temperature. We show how the whole set of experimental data supports an exciton self-trapping mechanism occurring at the interface of hBN-hBN homostructures.
@@
All CL experiments were performed using a JEOL7001F scanning electron microscope (SEM) equipped with a Horiba Jobin-Yvon CL system optimized for UV detection, as described in detail in reference [51]. Secondary electrons are detected by a Everhart-Thornley detector (SE detector). The SEM image corresponds to the intensity of the signal collected by the SE detector as a function of the position of the focused electron beam, with an ultimate spatial resolution of 1.2 nm. CL images are measured by a photomultiplier (Hamamatsu HJY model - cooled to -30°C by Peltier effect) as the electron beam scans the selected area of the sample. The signal from the SE detector and the intensity collected by the photomultiplier are recorded simultaneously. In this way, SEM and CL images of the same area are generated, which allows to correlate the emission of a CL signal with the topography of the sample. The CL image can be panchromatic (total intensity, unfiltered) or monochromatic, filtered in energy by a monochromator equiped with diffraction gratings. CL spectra are recorded with a CCD camera.
@@
Thanks to a careful calibration of the detection system, the CL signal provides the absolute intensity of light emission from the sample [51]. As a result, the internal quantum efficiency ($IQE$) of a luminescence signal can be evaluated from a CL experiment. The $IQE$ corresponds to the ratio between the rate of photons emitted inside the sample and the rate of electron-hole pairs generated by the electron bombardment. Details of the $IQE$ measurements for the considered signal are given in the Appendix A. Finally, TRCL experiments are performed using a custom-built beam-blanking device mounted on the SEM column. The overall time resolution of the TRCL set-up is measured equal to 100 ps (details of the setup are presented in ref. [50]).
@@
To study the emission of twisted hBN-hBN structures, we decided to vary a large number of parameters in the fabrication of the homostructures: the twist angle, the thickness of the crystal flakes in stack, but also the quality of the starting hBN crystals and the way the two crystal flakes are assembled. hBN crystals with natural boron isotope content were grown by three different processes: an Atmospheric Pressure High Temperature (APHT) process [52, 53, 54, 55] from Ni/Cr solvent, a High-Pressure High Temperature (HPHT) route [56, 57], and a Polymer Derived Ceramic (PDC) method [58, 59]. hBN from the HPHT method is recognized in the 2D materials scientific community as the reference hBN crystals. In a previous study, a quantitative benchmarking of the respective quality, in terms of defect density, of these different hBN sources was performed based on the measured free exciton lifetimes and correlated to the electron mobility in graphene/hBN h2Ds [50, 60]. The values of the lifetime and $IQE$ of the free exciton measured on bulk crystals from the different hBN sources are given in Tab. 1.
Sample Free exciton $\tau$ (ns) Top hBN Bottom hBN
$IQE$ (%) [50] [50] (nm) (nm)
@@
To fabricate twisted homostructures made of HPHT crystals, the bottom hBN layer is exfoliated with a PDMS stamp and deposited onto a Si/SiO2 substrate [61]. The second hBN flake is similarly exfoliated, and deposited over the bottom hBN under a microscope. The samples are then annealed at 150°C for 1.5 hours. With this method, one atomic layer of the twisted interface was in contact with the PDMS.
For twisted homostructures of APHT and PDC crystals, the bottom layer is peeled off using Scotch tape on a Si/SiO2 substrate [62], while the top layer is peeled off using PDMS [61] and deposited over the bottom hBN flake under the microscope. This method ensures that the two atomic layers of the hBN-hBN interface have never been in contact with either the PDMS or the Scotch tape and does not involve annealing. All the procedures, for HPHT, APHT and PDC samples were performed in a controlled atmosphere glove box.
The twist angle between the two hBN flakes is measured with diffraction by electron backscatter diffraction (EBSD) or by electron channeling patterns (ECP) as shown in Fig. 2. Details of the twist angle measurement are presented in the Appendix B. The samples are named according to the crystal synthesis method from which they are issued (HPHT, APHT or PDC) and the twist angle. Tab. 1 lists the different samples and the respective thickness of the top and bottom parts measured by Atomic Force Microscopy (AFM).
.3 III - Evidence of the excitonic origin of the broad 4-eV optical emission
@@
All hBN-hBN samples were first analyzed using continuous excitation CL at room temperature. The first observation is that the 4-eV luminescence mainly appears in homostructures with high twist angles (Fig. 3). At low twist, such as the PDC-2°, APHT-3° and APHT-4° samples, no clear luminescence signal could be detected in the 3-5 eV range, confirming a recent study [39]. We conclude that the 4-eV luminescence becomes truly significant at angles larger than 5°. A spectrum measured on a low angle hBN-hBN structure is presented in the Appendix C. Spectra measured on 3 highly twisted hBN-hBN homostructures (purple) and on a single hBN crystal as a reference (black) are shown in Fig. 3(b). The 4-eV emission is modulated by Fabry-Pérot interferences whose periodicity is consistent with the thickness of the flakes (see Tab. 1). CL images filtered at 4 eV are shown in Fig. 3(c). They clearly demonstrate that the 4 eV cathodoluminescence component is due to the presence of the twisted hBN-hBN interface.
@@
Figure 3: (a) SEM images of the twisted homostructures, outlined with a dotted line; (b) CL spectra measured on the bottom crystal alone (in black) and on the homostructure (in purple) at 300 K; (c) CL images filtered at 4.1$\pm$0.1 eV at 300 K, on HPHT-11° (5 kV, 27 pA), APHT-29° (3 kV, 280 pA), and PDC-18° (5 kV, 340 pA) samples. Scale bars are 10 µm.
The full set of spectra and CL images recorded for the 16 structures investigated in this work are provided in the Supplementary Information. The signals are comparable for all samples regardless of hBN source and flake thickness. They show common features in line with previous observations [39, 40, 41]: intense, 2-eV broadening and 2-eV redshift with respect to the 6.25-eV hBN bandgap [51].
In summary of these first observations, the emission phenomenon is very robust and appears unconditionally as soon as the twist angle is large, as already reported in Ref. [39] for instance. As it does not depend on the differences in nature and density of defects that present the three hBN sources used to fabricate our homostructures, the emission seems to be intrinsic to the presence of the interface itself rather than related to extrinsic defects. This raises the question of its origin, and more specifically, the interplay between the exciton and the twisted-interface structure, which we will explore next using TRCL.
@@
The goal of the TRCL study is to identify the role of bulk free excitons in the process leading to the 4-eV luminescence. Figure 4(a) compares the decay of the free exciton luminescence at 5.75 eV in the HPHT-11° homostructure with the decay obtained on the bottom hBN crystal alone. A drastic change is observed. For the crystal alone, the decay is dominated by a short 3 ns component, typical of the TRCL decay of the free exciton in a single bulk HPHT-hBN crystal [50]. For the homostructure, the fast initial component is similar but a long component of 0.46 µs appears with a significant intensity (30% on the HPHT-11° sample). This long time constant is much larger than the radiative lifetime of the free exciton measured at 27 ns [50], which requires an incoming flux of free excitons that persists after the primary excitation has stopped. We conclude that the long lasting free exciton emission is fed by the detrapping of excitons from the interface, revealing a balance between a population of interface excitons and the population of free excitons in the hBN volume based on a trapping-detrapping process.
@@
Figure 4: (a) Decay of the free exciton luminescence, filtered at 5.75$\pm$0.2 eV, on the bottom hBN alone (gray) and on the hBN-hBN twisted structure HPHT-11° (orange), after interrupting the TRCL excitation at initial time (5 kV, 27 pA, spot size of 10 µm). Bi-exponential decay functions are used for the fits. (b) Decay of the free exciton luminescence (black), and of the the 4-eV emission (purple), filtered at 4.1$\pm$0.1 eV, after interruption of steady excitation at t=0, on HPHT-11° homostructure (5 kV, 27 pA, 10 µm excitation spot diameter). Intensities during steady state excitation (t $<$ 0) are normalized by the integrated intensity of the bands recorded on continuous CL spectra. Experiments done at 300 K.
As shown in Appendix D, the temporal decay of the 4-eV luminescence is energy-independent in the 3-5 eV energy range. Figure 4(b) further illustrates that its decay time matches the long-lived component of the free exciton emission. Without going into the quantitative analysis (presented later), it demonstrates that the broad 4-eV luminescence is excited by the long-lived and interface-localized excitons. This implies that the 2 eV broadening likely does not arise from energy dispersion of an ensemble of localized quantum emitters. Instead, it appears to result from the radiative recombination process of interface excitons. Still, the radiative process has to be identified, which is the subject of the next parts.
.4 IV - Power dependence and internal quantum efficiency of the 4-eV emission.
@@
During CL experiments, the electron beam excitation is spread in depth, so that the hBN-hBN interface is indirectly excited by the diffusion of free excitons generated in the volume of the hBN crystals towards the interface between them [63]. The efficiency of the interface exciton luminescence, therefore, depends on two factors: the efficiency of free exciton transport from the bulk to the interface, and the efficiency of radiative recombination at the interface.
@@
The luminescence from the twisted hBN-hBN homostructures is analyzed here as a function of the areal density of the excitation power (W/cm2). Low power density data were recorded with an acceleration voltage of 5 kV with a constant current of 27 pA and spot diameters ranging from 1.6 to 26 µm. High power densities were obtained with a fixed spot diameter of 1.6 µm and excitation currents ranging from 27 to 5600 pA. This dual approach made it possible to vary the excitation power density over 5 orders of magnitude.